Lactic acid CH3CHOHCOOH is found in sour milk, in sauerkraut, and in muscles after activity. Lactic acid is a weak monoprotic acid in aqueous solution, with a Ka = 1.4x10^-4 at 25 degree. 2.15g of sodium hydroxide was added to 8.00x10^2 mL of an aqueous solution containing 0.100mol/L lactic acid.

a) Determine the pH of the resulting solution

b)determine the concentrations of all neutral acid ionic species (other than water) present in the resulting solution
Answer #1

pKa = 3.85

moles of NaOH = 2.15 / 40 = 0.05375 mol

moles of lactic acid = 0.100 x 0.800 = 0.0800 mol

a)

HA +    NaOH     ---------------->   NaA   +    H2O

0.08          0.05375                            0               0

0.02625        0                                0.05375

pH = pKa + log [salt / acid]

      = 3.85 + log [0.05375 / 0.02625]

pH = 4.16

b)

[Na+] = 0.05375 / 0.8

[Na+] = 0.0672 M

     =